OpenStudy (anonymous):

which line is parallel to the line segment joining P (-7,2) and Q (3, -8)

OpenStudy (anonymous):

first find the slope of a line passing through two given points? that is( -8 -2)/(3+7) =-10/10 =-1

OpenStudy (anonymous):

since parallel lines have same slope use this condition to find the equation of line parallel to it

OpenStudy (anonymous):

i have 4 options y=3/2x+2, b) y= -x+2 c) y= x+2 d) y= -3/2x+2 i think its c is it really?

OpenStudy (anonymous):